OpenStudy (anonymous):

Solve the inequality. -12 <= 2x - 4 < 10

mathslover (mathslover):

\[\huge{-12\le 2x-4 < 10}\] \[\huge{-12+4\le 2x-4+4 <10+4}\] I just added 4 in all inequalities \[\huge{-8\le 2x < 14}\] now divide all inequalities by 2 what you get ? may i know ?

OpenStudy (anonymous):

−8≤2x<14, expressed as absolute value.

OpenStudy (radar):

Come on, someone divide thru by 2.

OpenStudy (radar):

@rayjohnson . we are all waiting, with bated breath for the answer/.